OpenStudy (anonymous):

subtracting polynomials (7x+8)-(4x-4) =7x+8-(-4x)+4 =7x+8+4x+4 =11x+12?

OpenStudy (anonymous):

hey easy few steps

OpenStudy (anonymous):

u got 2 factors

OpenStudy (anonymous):

\[(7x+8)-(4x-4)\]

OpenStudy (anonymous):

now tell me what's \[+ \times -=?\] \[- \times - = ?\]

OpenStudy (anonymous):

+(-)=? -(-)=?

OpenStudy (anonymous):

+

OpenStudy (anonymous):

@juelz223

OpenStudy (anonymous):

+ for what?

OpenStudy (anonymous):

+(-) = - -(-) = +

OpenStudy (anonymous):

correct !!!!!!!!!!

OpenStudy (anonymous):

so since u know that u r closer to ur ans

OpenStudy (anonymous):

just open the brackets keeping the sign rules in mind take it easy

OpenStudy (anonymous):

i know man but i'm here on my pearsons math lab, doing homework and it's saying that 11x+12 is wrong

OpenStudy (skullpatrol):

How do you subtract real numbers?

OpenStudy (anonymous):

\[7x+8-4x+4\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

3x + 12

OpenStudy (skullpatrol):

Yes

OpenStudy (anonymous):

correct wow got it!!!

OpenStudy (anonymous):

idk what i did wrong there.. i mean this comes really simple to me and i got all the other ones right

OpenStudy (anonymous):

-(4x-4)=-4x+4

OpenStudy (anonymous):

got it? @juelz223

OpenStudy (skullpatrol):

To subtract a real number you add the opposite of that number.

OpenStudy (anonymous):

yeah you switch the signs of the subtracted number and add

OpenStudy (skullpatrol):

To subtract polynomials you add the opposite of each term and then simplify.

OpenStudy (anonymous):

so take care of that

OpenStudy (skullpatrol):